A quarter of a circle has a radius of 8 cm. Calculatexcept it's perimeter
1
Expert's answer
2020-02-24T11:31:33-0500

"P=\frac{2\pi r}{4}+2r=\frac{2\pi *8}{4}+2*8=16+4\pi\approx28.57 cm."